Let the six-month interest rate be 3% in UK, the spot rate be $1.60/₤, and the six-month forward rate be $1.61/₤. Assume the IRP condition holds. Find the six-month interest rate in the U.S.

please show steps
Answer #1

Forward Rate = Spot Rate * (1 +U SAInterest Rate/2) (1 +UKInterest Rate/2)

1.61 = 1.60 * (1 + USAInterest Rate/2) (1 +0.03/2)

1.61 / 1.60 = (1 + USA Interest Rate /2) / 1.015

1.00625 * 1.015 = 1 + USA Interest Rate /2

1.02134375 -1 = USA Interest Rate /2

0.02134375 * 2 = USA Interest Rate

0.0426875 = USA Interest Rate

USA Interest Rate = 4.27%

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.